A profit maximising firm in a competitive market is currently producing 50 units of output. The market price is Rs. 10, the firms average variable cost is Rs 5 and its fixed cost is Rs. 250.

(1) What is the profit of the firm? What is its marginal cost?

(iii) What is the efficient scale of production for this firm


1
Expert's answer
2021-03-03T07:54:13-0500
1. Profit = Revenue - Total cost
Revenue = Number of units sold multiplied by the average total cost
The average total cost = Total cost divided by the quantity
The total cost = fixed cost + variable cost
Thus Profit = (50*5.1 - 255) = 0
Marginal cost = change in the total cost divided by change in the quantity
Thus marginal cost = 255/50 = 5.1
ii)The efficient sclale = ratio of the marginal cost and the total average cost
Thus the efficient scale of this firm is 5.1/255 = 0.02